Using 1 variable like w=width than writing an inequality how would you work this problem?
The length of a rectangular boat storage shed must be 4 meters more
than the width. The perimeter must be at least 120 meters. What widths
of the shed will make the perimeter at least 120 meters?

Answer by rothauserc(4718)

we are given
The length of a rectangular boat storage shed must be 4 meters more
than the width. The perimeter must be at least 120 meters. What widths
of the shed will make the perimeter at least 120 meters?
w is width and l (length) = w + 4, also P (perimeter), we have
2w + 2*(w+4) = P > or = 120
2w + 2w + 8 > or = 120
4w > or = 112
w > or = 28 meters